### 29 Mar 2026 v2

# A Full-Density Approach to Simulating Random Iteration Equations with Applications

Abstract
The goal of this study is to introduce a unified computational framework for simulating random iteration equations (RIE), understood as iteration equations containing random variables. The novelty of this work is that full probability densities of the state vectors are propagated stepwise through the iterations avoiding the need of repetitive pathwise Monte Carlo simulations of the iteration equation. The presentation of the methodology is conceptually efficient based on recent work on static random equations and intentionally accessible. The technical requirements on the RIE are minimal based on the previous work, allowing for potential nonlinearities, discontinuities and stochasticities in the transfer function, as well as nonstandard densities and diffusion processes. As results, illustrative applications of random and stochastic differential equation simulations, a novel full-density gradient descent method (FDGD) for global optimization under uncertainty and examples of chaotic mappings are presented in order to demonstrate the breadth of the utility of this framework. In total, the character of the presentation is explorative and encourages new applications and theoretical studies.
Keywords: random iteration equation, density propagation, uncertainty quantification, stochastic dynamics, full-density gradient descent, stochastic differential equation, simulation
About the Author
Dr. Högele is Professor of Applied Mathematics and Computational Science at the Department of Computer Science and Mathematics at the Munich University of Applied Sciences HM, Germany. His research interests are mathematical and stochastic modeling, simulation and analysis of complex systems in applied mathematics with specific applications to radiotherapy, imaging, optical metrology and signal processing.
1 Introduction
In this work a novel computational framework is presented to investigate the evolution of random iteration equations (RIE). In general, first order iteration equations are transforming a state vector $\boldsymbol{x}\in\mathbb{R}^{R}$ starting from an initial state $\boldsymbol{x}^{(1)}$ with a stepwise iteration mapping from the $n$-th iteration state $\boldsymbol{x}^{(n)}$ to the $(n+1)$-th iteration state $\boldsymbol{x}^{(n+1)}$. This is denoted by the transfer function $\boldsymbol{T}:\mathbb{R}^{R}\rightarrow\mathbb{R}^{R}$ with $\boldsymbol{x}^{(n+1)}=\boldsymbol{T}(\boldsymbol{x}^{(n)})$. RIEs will be defined as iteration equations that contain random variables, i.e. $\boldsymbol{T}$ contains stochasticities summarized by random variables in the iteration equation and, in consequence, the resulting next state $\boldsymbol{x}^{(n+1)}$ must also be interpreted as a random variable with a probability density. This means, the RIE gets the density function of a previous state and computes the density function of the next state by incorporating the random variable densities in the transfer function. In total, we are observing the stepwise evolution of state probability density functions from iteration to iteration, being an example of a density propagation algorithm [1, 2]. The proposed framework shows how these density functions can be computed in practice which is demonstrated by expressive examples. RIEs can be regarded as a key component for the simulation of many dynamical systems that transform density functions, which places this study in the broad field of uncertainty quantification [3]. A minor point in this argumentation, but a major difference in the application, is that also the initial state $\boldsymbol{x}^{(1)}$ will be treated as a random variable, i.e. the iteration begins not with single starting values, but with a full distribution of them.
In order to understand the broader relevance of having such a computational framework, this stochastic extension is presented for three major application areas of iteration equations in applied mathematics: simulation of dynamical systems (including random and stochastic differential equations), optimization (extending gradient descent optimization to a full-density gradient descent method) and chaotic maps (presenting the strange attractors of the Ikeda map and the Lozi map). In no way this is the full potential of this framework, instead it just shows its relevance for undeniably important areas in applied mathematics utilizing expressive example simulations.
Standard approaches to simulate dynamics depending on uncertainties or diffusion processes are performing statistical pathwise Monte Carlo simulation, i.e. from all probability densities a single sample is drawn and the corresponding deterministic iteration is performed. This pathwise approach is performed in repetition and all results are collected statistically in a histogram map in order to see the final density of paths [4, 5]. This is not an evolution or propagation of probability densities, but a purely statistical approach. Based on a recent publication by Hoegele [6] which focused on a method to compute the probabilistic solution space of random equations, the propagation of the probability densities can be performed directly by computing the full state density from iteration to iteration considering the complete stochastic environment. Although in the numerical framework we utilize Monte Carlo integration techniques, this should not be misinterpreted with the statistical collection of pathwise Monte Carlo simulations.
It is clearly stated, that this paper presents a simulation framework as mentioned prominently in the title with explorative and expressive simulation examples in order to present the capabilities of this computational framework. It is not deriving limit theorems, convergence rates and theoretical bounds. Major contributions on this theory is done with the closely related concept of iterated random functions which are studied mainly for the existence of stationary distributions utilizing Lipschitz contraction arguments [7, 8], while our approach is focusing on the transient phase and explicit evolution of the density propagation through stepwise nonlinear iterations. Specifically, this study demonstrates that the density propagation of stochastic and dynamic iteration problems can be simulated straightforwardly, even if the theoretical treatment might be enormously difficult or convergence is not guaranteed or (partially) violated. In our perspective, this is more aligned with real world applications in which mostly only a limited number of iterations are of computational interest. Moreover, we typically utilize normal distributions for the parameter random variables or diffusion processes, but this is not a limitation of the computational framework and non-standard probability densities (also non-standard diffusion processes) combined with significant nonlinearities or discontinuities or without Lipschitz contraction properties of the transfer mapping are as easily handled within this framework as more well-studied situations.
For example, specifically with respect to stochastic differential equations this approach is therefore a computational complement to Perron-Frobenius-Operator approaches (which are typically applied to deterministic, non-random transfer functions and focus on the limiting density) and Fokker-Planck-Equation approaches (which are difficult to apply if continuous-time Gaussian diffusion processes or smooth drift or diffusion coefficients are not guaranteed). And on the computational side it is also complementary to pathwise Monte Carlo simulations due to the focus on full-density propagation.
2 Methods
2.1 General Formulation
We introduce the random iterative equation (RIE) with a (typically at least) piecewise continuous transfer function $\boldsymbol{T}:\mathbb{R}^{R}\times\mathbb{R}^{K}\rightarrow\mathbb{R}^{R}$ and a random variable parameter vector $\boldsymbol{C}^{(n)}$ with $K$ entries and known densities (depending on the iteration number $n$) by
$\boldsymbol{x}^{(n+1)}=\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})\;,$ (1)
given an initial state $\boldsymbol{x}^{(1)}\sim f_{\boldsymbol{x}^{(1)}}$. We want to utilize the methodology recently introduced by Hoegele [6] to rewrite this into the standard random equation form $\boldsymbol{M}(\boldsymbol{x};\boldsymbol{A})=\boldsymbol{B}$ with independent random variable vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, and $\boldsymbol{x}$ the corresponding solution vector. We transform the iteration equation
$\boldsymbol{x}^{(n+1)}-\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})=\boldsymbol{0}$ (2)
leading to $\boldsymbol{B}$ a random variable vector centered around zero and with small standard deviations compared to other standard deviations and the required solution vector $\boldsymbol{x}^{(n+1)}$ for each iteration. The rather technical minimal assumptions about $\boldsymbol{T}$ are inherited directly from the integrability assumptions about $\boldsymbol{M}$ presented in [6] and are typically only violated in constructed limit cases. In total, this leads to
$\boldsymbol{M}:\mathbb{R}^{R}\times\mathbb{R}^{R+K}\rightarrow\mathbb{R}^{R}$ (3)
$\boldsymbol{M}(\boldsymbol{x}^{(n+1)};\underbrace{\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)}}_{=:\boldsymbol{A}}):=\boldsymbol{x}^{(n+1)}-\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ (4)
with $\boldsymbol{A}:=(x_{1}^{(n)},\dots,x_{R}^{(n)},C_{1}^{(n)},\dots,C_{K}^{(n)})^{T}$ the random variable vector consisting of two parts where $\boldsymbol{x}^{(n)}$ is a $R$-dimensional random vector with density $f_{\boldsymbol{x}^{(n)}}$ computed in the $n$-th iteration, and $\boldsymbol{C}^{(n)}$ a static $K$-dimensional random vector with i.i.d. $C_{1}^{(1)},\dots,C_{1}^{(n)}\sim f_{C_{1}}$ for the first entry, until i.i.d $C_{K}^{(1)},\dots,C_{K}^{(n)}\sim f_{C_{K}}$ for the last entry. The common density function is denoted by $f_{\boldsymbol{C}}$. For a well-defined iteration a starting density $f_{\boldsymbol{x}^{(1)}}$ must be defined. The independency of the $C_{i}^{(n)}$ for different $n$ is directly in the spirit of deterministic iteration equations, where the transfer function itself is independent of previous evaluations of it.
Since $\boldsymbol{x}^{(n)}$ and $\boldsymbol{C}^{(n)}$ are independent, we get a likelihood distribution for $\boldsymbol{x}^{(n+1)}$ based on those densities and the specific transfer function $\boldsymbol{T}$ by
$\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})=\int\limits_{\mathbb{R}^{K-R}}\int\limits_{\mathbb{R}^{R}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\;\text{d}\boldsymbol{s}_{1}\,\text{d}\boldsymbol{s}_{2}\;.$ (5)
which general form was presented previously [6]. When focusing on probability densities, we reformulate the likelihood function with Bayes’ theorem to a posterior density by utilizing a non- or weak-informative prior in this study which does not interact with the likelihood intensities by
$\pi_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{x})\propto\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})\;,$ (6)
essentially normalizing the likelihood function to an integral value of $1$. In consequence, we can calculate in each iteration the posterior density function $f_{\boldsymbol{x}^{(n+1)}}:=\pi_{\boldsymbol{x}^{(n+1)}}$, evolving from a prior starting density $f_{\boldsymbol{x}^{(1)}}:=\pi_{\boldsymbol{x}^{(1)}}$. Of course, in each iteration there could also be a density modifying prior utilized which strongly influences the density propagation. This could be important, for example, if the state space is restricted by hard or soft boundaries (utilizing a uniform prior in the region of admissible solutions) and/or is a discrete space with discrete probability densities (utilizing a gridded prior). We will leave such interesting possibilities outside of this study (although it is not outside of the framework) and only refer to the corresponding Section 2.5 in Hoegele [6].
2.2 Numerical Implementation
The main goal of this work is to present a computational framework for calculating the iterations of the RIE. As presented by Hoegele [6], the integral in Equation 5, can be approximated by Monte Carlo integration, utilizing $P$ samples with the formula
$\pi_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{x})\propto\;\frac{1}{P}\sum\limits_{p=1}^{P}\left(\prod\limits_{r=1}^{R}f_{B_{r}}(x_{r}-T_{r}(\boldsymbol{s}_{1,p},\boldsymbol{s}_{2,p}))\right)\;,$ (7)
where the samples $\boldsymbol{s}_{1,p}$ are drawn form the density $\pi_{\boldsymbol{x}^{(n)}}$ and $\boldsymbol{s}_{2,p}$ are drawn from $f_{\boldsymbol{C}^{(n)}}$. The sampling for the latter is straightforward, since these are given by the RIE definition and one can for example utilize latin hypercube sampling (even the same representative samples for all iterations once calculated before iterations start). The sampling of the first density $\pi_{\boldsymbol{x}^{(n)}}$ must be done in each iteration and these densities are non-trivial. There are standard Monte Carlo Markov Chain (MCMC) methods in order to draw from such posteriors, but in this work we prefer due to simplicity an acceptance rejection approach [4]. Essentially, drawing in a $R$-dimensional rectangular box around the density region of $\pi_{\boldsymbol{x}^{(n)}}(\boldsymbol{x})$ uniformly distributed samples $\boldsymbol{x}_{p}$ with one additional uniform sample $y_{p}$ in the range of 0 and the maximum overall value of $\pi_{\boldsymbol{x}^{(n)}}$. If $y_{p}$ is larger than $\pi_{\boldsymbol{x}^{(n)}}(\boldsymbol{x}_{p})$ it is rejected, else it is kept. This is done in parallelization for speed-up and exactly $P$ samples are computed in each iteration.
A possible schematic algorithm (which is also the basis of the presented results) looks like
  • 1.
    Define the RIE, especially the transfer function $\boldsymbol{T}:\mathbb{R}^{R}\times\mathbb{R}^{K}\rightarrow\mathbb{R}^{R}$, the i.i.d. random variable densities $f_{\boldsymbol{C}}$ and the initial prior density $\pi_{\boldsymbol{x}^{(1)}}$. Set $n=1$.
  • 2.
    Take $P$ draws $\boldsymbol{s}_{2,p}$ ($p=1,\dots,P$) from $f_{\boldsymbol{C}}$ by standard sampling methods (e.g. by latin hypercube sampling).
  • 3.
    Take $P$ draws $\boldsymbol{s}_{1,p}$ ($p=1,\dots,P$) from $\pi_{\boldsymbol{x}^{(n)}}$ (e.g. by acceptance rejection).
  • 4.
    Evaluate Equation 7 numerically in order to calculate $\pi_{\boldsymbol{x}^{(n+1)}}$ (including the normalization step of the approximated density) on the same grid as $\pi_{\boldsymbol{x}^{(n)}}$ was provided.
  • 5.
    Set $n\rightarrow n+1$ and repeat steps 3–5 until a certain limit has been reached, such as a given iteration number or the changes between $\pi_{\boldsymbol{x}^{(n)}}$ and $\pi_{\boldsymbol{x}^{(n+1)}}$ are below a certain tolerance measure.
This algorithm for simulating full densities of the RIEs can be applied directly for the stochastic extension of a given iteration equation, such as the chaotic mappings presented in the results Section 3.3. More important in applied mathematics are situations where the iteration equation itself (and its stochastic extension to RIE) is a tool to solve challenging (mostly nonlinear) problems. We will present two of the most prominent use cases in the next subsections.
2.3 Discretization of Dynamical Systems with Stochasticity as Random Iteration Equation
We will see, that being able to simulate RIEs are a key step to simulate approximately random differential equation (RDE) and stochastic differential equation (SDE) dynamical systems in order to observe their evolution of the full densities over time [9, 5]. We are starting applications with a general ODE system with random variable parameters $\boldsymbol{C}$, representing a RDE,
$\dot{\boldsymbol{x}}=\boldsymbol{F}(t,\boldsymbol{x},\boldsymbol{C})\;,$ (8)
with a well-defined right hand side $\boldsymbol{F}$. By applying the discretization $\dot{\boldsymbol{x}}\approx\frac{\boldsymbol{x}^{(n+1)}-\boldsymbol{x}^{(n)}}{\Delta t}$, this leads to the iteration equation
$\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}+\Delta t\cdot\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})\;,$ (9)
essentially representing a random Euler iteration equation. When using the nomenclature of Section 2.1, this leads to the general transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)}):=\boldsymbol{x}^{(n)}+\Delta t\cdot\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})\;,$ (10)
for an arbitrary dynamical system. Of course, more advanced discretization schemes can be utilized, but for simplicity we leave this demonstration applying the Euler algorithm.
A major observation is that when utilizing this discretization scheme, we can also simulate the evolution of probability density functions of SDEs with random coefficients utilizing the Euler-Maruyama algorithm [4, 5, 10] by introducing new terms
$\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}+\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})\,\Delta t+\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{D}^{(n)})\,\Delta\boldsymbol{W}^{(n)}\;,$ (11)
with the diffusion coefficient function $\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{D}^{(n)})$, which may depend on a random variable vector $\boldsymbol{D}^{(n)}$, and a Wiener process $\boldsymbol{W}_{t}$ with a discrete increment $\Delta\boldsymbol{W}^{(n)}=\boldsymbol{W}_{t_{n+1}}-\boldsymbol{W}_{t_{n}}$ as normally distributed random variables with zero mean and variances $\Delta t$. This leads to the general transfer function for SDEs
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)}):=\boldsymbol{x}^{(n)}+\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}_{1..K_{1}}^{(n)})\,\Delta t+\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}_{K_{1}+1..K_{1}+K_{2}}^{(n)})\,\boldsymbol{C}_{K_{1}+K_{2}+1..K_{1}+K_{2}+R}^{(n)}\;,$ (12)
with three parts of $\boldsymbol{C}$, i.e. the $K_{1}$ random variables of the drift coefficient $\boldsymbol{C}_{1..K_{1}}$, the $K_{2}$ random variables of the diffusion coefficient $\boldsymbol{C}_{K_{1}+1..K_{1}+K_{2}}$ and the $R$ random variables of the discrete increment in the Wiener process $\boldsymbol{C}_{K_{1}+K_{2}+1..K_{1}+K_{2}+R}$.
Please note, that Equation 12 is a general model including parameter uncertainties (the RDE part) as well as a diffusion process (the SDE part) in one single computational framework. This shows already from the methodological side the power of the RIE approach for approximating those in general challenging processes combined. Additionally, it can be observed that if $\boldsymbol{C}^{(n)}$ (and respectively $\boldsymbol{D}^{(n)}$) of Equation 11 are independent (but not necessarily identically distributed) for different $n$ then also more complicated coefficient processes can be simulated just by reinterpretation of the terms (not leaving the computational framework). This exemplifies the conceptual similarities between RDEs and SDEs, especially from a simulation perspective. Further, the SDEs are not limited to Wiener processes, but can in the simulation take any other stochastic process that increment can be calculated for each iteration (which is deliberately general and leaves a lot of possibilities). Finally, also more sophisticated approaches than the simplistic Euler-Maruyama algorithm can be utilized for SDE simulation, without leaving the presented general framework as long as they lead to iteration equations. For the convergence rates of standard SDE types for the Euler-Maruyama algorithm see e.g. Higham[10].
Obviously, this strict focus on the discretized iteration equation as the key quantity opens up many possibilities for simulating challenging RDEs and SDEs besides the standard theory and current research in their fields. In this study, this is only regarded as one example application of RIE simulations.
2.4 Full-Density Gradient Descent as Random Iteration Equation
The idea of gradient descent is to update a current position $\boldsymbol{x}^{(n)}$ in a minimization process of an objective function $F(\boldsymbol{x})$ with $F:\mathbb{R}^{R}\rightarrow\mathbb{R}$ stepwise in the direction of the steepest downward direction which is represented by the negative direction of the local gradient, i.e. leading to the gradient descent (GD) algorithm
$\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-\eta\cdot\nabla F(\boldsymbol{x}^{(n)})\;,$ (13)
with a learning rate $\eta$. We present several stochastic extensions to this classical optimization routine, which family will be denoted with full-density gradient descent (FDGD):
  • 1.
    FDGD-I: Minimization of a stochastic objective function containing $K$ random variables for parameters of the objective function
    $\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-\eta\cdot\nabla F(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ (14)
    In this version stochastic optimization is performed, i.e. we have an objective function containing random variable parameters. We just want to present this approach here since it is an obvious version, but are not studying it in detail further on.
  • 2.
    FDGD-II: Minimization of a deterministic objective function updating the deterministic direction simultaneously for a full density of directions, leading to $R$ random variables
    $\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-\eta\cdot\nabla F(\boldsymbol{x}^{(n)})+\boldsymbol{C}^{(n)}$ (15)
    This approach introduces a diffusion process into the iteration which helps exploring the objective function.
  • 3.
    FDGD-III: Additionally to FDGD-II the update utilizes a full density of learning rates, leading to $R+1$ random variables
    $\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-C_{R+1}^{(n)}\cdot\nabla F(\boldsymbol{x}^{(n)})+\boldsymbol{C}^{(n)}\;,$ (16)
    leading to an update with different optimization schemes simultaneously.
FDGD-II and FDGD-III contain a diffusion component due to the added random variables in each iteration with a simultaneous convergence to local minima, producing complex dynamics during optimization. This leads, for example, to the transfer function for FDGD-III
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)}):=\boldsymbol{x}^{(n)}-C_{R+1}^{(n)}\cdot\nabla F(\boldsymbol{x}^{(n)})+\boldsymbol{C}^{(n)}\;,$ (17)
and for the other FDGD schemes analog. Again, this formulation is very general for any differentiable objective function $F$. In addition, combining FDGD-I and FDGD-III addresses an extremely broad set of stochastic optimization problems, transforming the local gradient descent algorithm to a global stochastic optimization technique. The difference to stochastic gradient descent (SGD) in the literature is that we do not use a single realization of the stochastic path in each iteration (which in SGD is produced by subsets of data for calculating the gradient in machine learning) but the whole density functions are transformed which allows a much broader search [11]. Related ideas of density propagation by stochastic gradient descent [12] and for the connection of SDEs and gradient descent [13] are part of recent research. We further want to draw the attention to the strong similarities of SDEs and the FDGD approach within the RIE framework which may lead to further insights.
3 Results
3.1 Evolution of Dynamical Systems
We present two versions of one dynamical system example. First, the normalized Rosenzweig McArthur as a RDE model [14, 15, 16] for predator prey simulations is given by
$\dot{x}_{1}$ $=x_{1}\,\left(1-\frac{x_{1}}{C_{1}}\right)-\frac{C_{2}\,x_{1}\,x_{2}}{1+x_{1}}$ (18)
$\dot{x}_{2}$ $=-C_{3}\,x_{2}+\frac{C_{2}\,x_{1}\,x_{2}}{1+x_{1}}\;,$ (19)
with the random variables representing uncertainties about the parameters of the model. The straightforward discretization $\dot{x}\approx\frac{x^{(n+1)}-x^{(n)}}{\Delta t}$ leads to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ $:=\left(\begin{array}[]{c}x_{1}^{(n)}+\Delta t\cdot\left(x_{1}^{(n)}\,\left(1-\frac{x_{1}^{(n)}}{C_{1}^{(n)}}\right)-\frac{C_{2}^{(n)}\,x_{1}^{(n)}\,x_{2}^{(n)}}{1+x_{1}^{(n)}}\right)\\ x_{2}^{(n)}+\Delta t\cdot\left(-C_{3}^{(n)}\,x_{2}^{(n)}+\frac{C_{2}^{(n)}\,x_{1}^{(n)}\,x_{2}^{(n)}}{1+x_{1}^{(n)}}\right)\end{array}\right)$ (22)
In the simulation, we utilize independent Gaussian densities $f_{C_{1}}:=\mathcal{N}(1,0.01^{2})$, $f_{C_{2}}:=\mathcal{N}(1,0.01^{2})$ and $f_{C_{3}}:=\mathcal{N}(0.25,0.01^{2})$ as well as $f_{B_{1}},f_{B_{2}}:=\mathcal{N}(0,0.005^{2})$. The differential equation model is known to lead to a steady state density [16] and we want to observe the evolution of the starting distribution to this steady state distribution based on the discretized iteration equation. We utilize $P=192000$ samples and $\Delta t=0.2$ (observing acceptable convergence results with minor numerical artifacts, see Figure 3 (left) for an impression). In Figure 1, we are utilizing starting density $f_{\boldsymbol{x}^{(1)}}$ as a uniform density on $0.1,0.5]\times[0.1,0.5]$ and observe its evolution into the steady state distribution. The deterministic steady state (when taking the modes of each density function as parameter values) is at $(\frac{1}{3},\frac{8}{9})$ and coincides well with the simulation result.
Figure 1: Illustration of the posterior densities for different iteration numbers of the Rosenzweig McArthur RDE model. One can see the evolution of the square uniform initial density and how it evolves due to the dynamics of the Rosenberg McArthur model with random variable parameters to the steady state density.
In the second example, we present a SDE version of the normalized Rosenzweig McArthur model [19], i.e.
$\text{d}x_{1}$ $=\left(x_{1}\,\left(1-\frac{x_{1}}{k}\right)-\frac{m\,x_{1}\,x_{2}}{1+x_{1}}\right)\,\text{d}t+\sigma_{1}\,x_{1}\,\text{d}W_{1}$ (23)
$\text{d}x_{2}$ $=\left(-c\,x_{2}+\frac{m\,x_{1}\,x_{2}}{1+x_{1}}\right)\,\text{d}t+\sigma_{2}\,x_{2}\,\text{d}W_{2}\;,$ (24)
with the real parameters $k,m,c$ and a Wiener process $\boldsymbol{W}_{t}$, leading to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ $:=\left(\begin{array}[]{c}x_{1}^{(n)}+\left(x_{1}^{(n)}\,\left(1-\frac{x_{1}^{(n)}}{k}\right)-\frac{m\,x_{1}^{(n)}\,x_{2}^{(n)}}{1+x_{1}^{(n)}}\right)\,\Delta t+\sigma_{1}\,x_{1}^{(n)}\,C_{1}^{(n)}\\ x_{2}^{(n)}+\left(-c\,x_{2}^{(n)}+\frac{m\,x_{1}^{(n)}\,x_{2}^{(n)}}{1+x_{1}^{(n)}}\right)\,\Delta t+\sigma_{2}\,x_{2}^{(n)}\,C_{2}^{(n)}\end{array}\right)\;.$ (27)
In this simulation we utilize a parameter set which leads to periodical orbits in the $x_{1}$-$x_{2}$-space in the deterministic case with $k=1.9,m=1.1,c=0.31$, but add the stochastic noise densities $f_{C_{1}},f_{C_{2}}:=\mathcal{N}(0,\Delta t)$ of the discritized SDE and use scaling coefficients $\sigma_{1}=\sigma_{2}=0.04$. In Figure 2 we utilize $P=384000$ samples and $\Delta t=0.05$ (leading to the observation of a clear orbital structure with minor numerical artifacts, see Figure 3 (right) for an impression) and present in the plot representative density functions (iteration 476 to 780) roughly over one period of the orbit. The periodicity is not perfect since the parameter values are a compromise for the periodicity of the full starting value range (which lead also to slightly different periodicities, see Figure 3 (right)) and the SDE diffusion leads to broad distributions of the orbital state density. Still, a clear orbit is observable also in this SDE version and it is demonstrated that such challenging simulations can be performed based purely on the computational framework of iterative density transformations.
An additional verification of the full-density simulation is provided in Appendix Section 6.1 where the simulation results are compared to the evolution of the analytically known mean and covariances of a 2D Ornstein-Uhlenbeck type SDE.
Figure 2: Illustration of the posterior densities for different iteration numbers of the Rosenzweig McArthur SDE model. The evolution of the densities is presented roughly over one period of the orbit in the $x_{1}$-$x_{2}$-space starting at iteration 476 ($t=23.8$), with dominant stochastic effects due to the SDE diffusion and initial state density.
Figure 3: Numerical impression of step sizes for the discretization of the dynamic systems by deterministic simulations in the relevant time and initial value range (two starting value examples, which are vertices of each initial density box) with ode45 (Runge-Kutta-(4,5) algorithm) as ground truth and the Euler algorithm which is basis of the simulaion strategy. Left: for system 1 utilized in the simulation in Figure 1 with starting values $(0.1,0.1)$ and $(0.5,0.5)$, right: for system 2 utilized in the simulation in Figure 2 with starting values $(0.4,0.4)$ and $(0.6,0.6)$. Sufficient numerical approximation accuracy in the context of this study can be observed considering the simulated stochasticities.
3.2 Optimization with Full-Density Gradient Descent
We present two expressive objective functions in order to demonstrate the use of this methodology for gradient descent. First, we present an objective function with two local minima and one saddle point in between
$F(\boldsymbol{x})=x_{1}^{4}-3\,x_{1}^{2}+x_{1}+5+2\,x_{2}^{2}\;,$ (28)
presented in Figure 4 on the left side with
$\nabla F(\boldsymbol{x})=\left(\begin{array}[]{c}4\,x_{1}^{3}-6\,x_{1}+1\\ 4\,x_{2}\end{array}\right)\;,$ (31)
and apply the FDGD-II algorithm, leading to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ $:=\left(\begin{array}[]{c}x_{1}^{(n)}-\eta\cdot\left(4\,\left(x_{1}^{(n)}\right)^{3}-6\,x_{1}^{(n)}+1\right)+C_{1}^{(n)}\\ x_{1}^{(n)}-\eta\cdot\left(4\,x_{2}^{(n)}\right)+C_{2}^{(n)}\end{array}\right)\;.$ (34)
Second, we present the famously challenging Himmelblau’s objective function [18, 13] with four minima which are difficult to localize
$F(\boldsymbol{x})=(x_{1}^{2}+x_{2}-11)^{2}+(x_{1}+x_{2}^{2}-7)^{2}\;,$ (35)
presented in Figure 4 on the right side with
$\nabla F(\boldsymbol{x})=\left(\begin{array}[]{c}4\,x_{1}\,(x_{1}^{2}+x_{2}-11)+2\,(x_{1}+x_{2}^{2}-7)\\ 2\,(x_{1}^{2}+x_{2}-11)+4\,x_{2}\,(x_{1}+x_{2}^{2}-7)\end{array}\right)\;,$ (38)
and apply the FDGD-III algorithm, leading to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ $:=\left(\begin{array}[]{c}x_{1}^{(n)}-C_{3}^{(n)}\cdot\left(4\,x_{1}^{(n)}\,\left(\left(x_{1}^{(n)}\right)^{2}+x_{2}^{(n)}-11\right)+2\,\left(x_{1}^{(n)}+\left(x_{2}^{(n)}\right)^{2}-7\right)\right)+C_{1}^{(n)}\\ x_{1}^{(n)}-C_{3}^{(n)}\cdot\left(2\,\left(\left(x_{1}^{(n)}\right)^{2}+x_{2}^{(n)}-11\right)+4\,x_{2}^{(n)}\,\left(x_{1}^{(n)}+\left(x_{2}^{(n)}\right)^{2}-7\right)\right)+C_{2}^{(n)}\end{array}\right)\;.$ (41)
In Figure 5, for the first objective function (Figure 4, left) we utilize independent Gaussian densities $f_{C_{1}}:=\mathcal{N}(0,0.04^{2})$ and $f_{C_{2}}:=\mathcal{N}(0,0.04^{2})$ as well as $f_{B_{1}},f_{B_{2}}:=\mathcal{N}(0,0.02^{2})$ and $P=48000$ samples. This FDGD-II example with learning rate $\eta=0.075$ with a small starting density shows how the gradient descent is attracted by the saddle point first and afterwards converges simultaneously to the two minima. In Figure 6, for a FDGD-III example with the second objective function (Figure 4, right) we utilize $f_{C_{1}}:=\mathcal{N}(0,0.2^{2})$, $f_{C_{2}}:=\mathcal{N}(0,0.2^{2})$ as well as $f_{B_{1}},f_{B_{2}}:=\mathcal{N}(0,0.02^{2})$ and the learning rate density $f_{C_{3}}:=\mathcal{N}(0.01,0.003^{2})$ (using $P=384000$ samples). Starting with a broad density, we see the fast attraction by a broad search for the posterior to the four minima. This shows, that in this stochastic regime an unspecific search strategy can be beneficial finding challenging extrema in a simultaneous fashion.
Figure 4: Illustration of the objective function. Left: With two local minima on the $x_{1}$-axis. Right: Himmelblau’s function of optimization for four local minima.
Figure 5: Illustration of the posterior densities for different iteration numbers for FDGD-II for the objective function with two local minima. One can see the evolution of the square initial density being first attracted to the saddle point and second the convergence to the two local minima.
Figure 6: Illustration of the posterior densities for different iteration numbers for FDGD-III using Himmelblau’s test objective function. One can see the evolution of the broad circular initial density with a variation of the update rate and its fast convergence to the four local minima.
3.3 Chaotic Random Iterations
We again present two numerical examples for this class of applications. Since chaotic random iterations are already applicable in the standard formulation of Section 2.1, we use the presented framework directly. First, the classical Ikeda map [19] in real space is presented by
$x_{1}^{(n+1)}$ $=1+u\cdot(x_{1}^{(n)}\cdot\cos(t^{(n)})-x_{2}^{(n)}\cdot\sin(t^{(n)}))$ (42)
$x_{2}^{(n+1)}$ $=u\cdot(x_{1}^{(n)}\cdot\sin(t^{(n)})+x_{2}^{(n)}\cdot\cos(t^{(n)}))\;,$ (43)
with
$t^{(n)}=0.4-\frac{6}{1+\left(x_{1}^{(n)}\right)^{2}+\left(x_{2}^{(n)}\right)^{2}}\;.$ (44)
Depending on the parameter $u$ different strange attractors appears. We are exploring the RIE for a small range of the variable $C_{1}^{(n)}=u$ represented by a random variable leading to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},C_{1}^{(n)})$ $:=\left(\begin{array}[]{c}1+C_{1}^{(n)}\cdot(x_{1}^{(n)}\cdot\cos(t^{(n)})-x_{2}^{(n)}\cdot\sin(t^{(n)}))\\ C_{1}^{(n)}\cdot(x_{1}^{(n)}\cdot\sin(t^{(n)})+x_{2}^{(n)}\cdot\cos(t^{(n)}))\end{array}\right)\;.$ (47)
In Figure 7 we explore the progress of iterations for $f_{C_{1}}:=\mathcal{N}(0.7,0.02^{2})$ with $P=384000$. It can be observed that the strange attractor arises during iterations, but also how the Ikeda mapping behaves in its first iterations acting on a full starting density distribution. This is an average of the strange attractor according to the ensemble distribution of the $u$-parameter ($C_{1}^{(n)}$).
Figure 7: Illustration of the posterior densities of the Ikeda mapping. One can see the evolution of the broad square initial density into the characteristic strange attractor of the Ikeda mapping averaged over a density of the $u$-parameter.
Second, the classical Lozi map [20] is presented, which is given by
$x_{1}^{(n+1)}$ $=1-a\cdot|x_{1}^{(n)}|+x_{2}^{(n)}$ (48)
$x_{2}^{(n+1)}$ $=b\cdot x_{1}^{(n)}\;,$ (49)
with typically a fixed standard parameter $b=0.3$ and a range for $a\in[1.4,1.7]$ for which strange attractors appear. This can be transferred to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},C_{1}^{(n)})$ $:=\left(\begin{array}[]{c}1-C_{1}^{(n)}\cdot|x_{1}^{(n)}|+x_{2}^{(n)}\\ b\cdot x_{1}^{(n)}\end{array}\right)\;,$ (52)
being an example for a non-differentiable transfer function. In Figure 8 we explore the progress of iterations for $f_{C_{1}}:=\mathcal{N}(1.55,0.1^{2})$ with $P=384000$. Also in this case the characteristic strange attractor with its sharp edges can be observed, even if we utilize a full ensemble of $a$ parameters.
Figure 8: Illustration of the posterior densities of the Lozi mapping. One can see the evolution of the broad square initial density into the strange attractor of the Lozi mapping with sharps edges averaged over a density of the $a$-parameter.
4 Discussion
In this explorative study we investigated a novel computational framework for RIE. This work is mainly based on the paper by Hoegele [6], which demonstrated how to calculate the probability density of the solutions space of (static and nonlinear) random equation systems. The main idea in this paper is now to utilize this methodology in an iterative fashion, leading to dynamics of probability density propagations. Essentially, the iteration transforms in each step the starting density function combined with the random variables contained in the iteration function itself in order to compute the next iteration state density. This can be regarded as computing ensemble solutions for dynamical systems, e.g. [21].
A major difference compared to pathwise Monte Carlo simulations is that the full density of the state space is calculated in each iteration step instead of several deterministic paths which are averaged afterwards. An advantage could be the adaptive and more efficient sampling of the occurring densities during simulation, since in the pathwise approach only after the full simulations the densities appear statistically.
As all iterative schemes, RIEs show interesting properties. This is demonstrated by the application to stochastic dynamical systems, such as RDEs and SDEs, which show nontrivial evolutions of the state densities. We also show how this framework can be used for an explorative gradient descent, which we denote by FDGD (full-density gradient descent) as a powerful extension of classical gradient descent to global optimization. Essentially, this leads to a meaningful search strategy for many local minima simultaneously with decisions taken not by single random paths but for a whole density of search directions. In this context, stochastic gradient descent [11] could be associated with single random path of optimization and ensemble optimization with the tracking of a number of sample paths in parallel during iterations, but not the full density. We want to emphasize that the presented simulation of dynamical systems and optimization are two main applications of applied mathematics, and utilizing RIEs allows a straightforward extension to stochastic regimes. Finally, we provide an impression how strange attractors in chaotic mappings are computed with this framework for full densities of starting values and parameters. This is the most direct application of the framework.
Another important task in applied mathematics is solving nonlinear equation systems, e.g. by Newton’s method or similar approaches, which typically lead in the deterministic case also to iteration equations. Since solving static equations (without a dynamic component) was already completely addressed in Hoegele [6] by the general nonlinear form $\boldsymbol{M}(\boldsymbol{x};\boldsymbol{A})=\boldsymbol{B}$ without the need of iteration equations, the extension to this computational framework is not necessary and it is referred to the paper.
The most important point of the presentation of this framework is that it is simple (if there is a method to solve static random equations) and very general, and it should be applicable to many dynamics which are based on iterations (explicitly focusing on discretized continuous phenomena). As a side note, in dynamical systems the time step $\Delta t$ could also be taken as random variable (similar to the learning rate in FDGD-III), representing different discretization schemes of the dynamical system simultaneously. This leads to a new understanding of the simulation of dynamical systems, allowing to stochastically explore the discretization method itself. In general, combining the two areas, discretized dynamical systems and gradient optimization within the RIE framework may lead to new insights in both areas.
This study is explorative and not a derivation of theoretical bounds or convergence rates of numerical schemes, which are important on their own and complementary to this presentation. Instead this study explores the possibility to approximately compute and visualize the evolution of a finite number of iterations. We regard this focus as underrepresented in the study of stochastic dynamical systems which is a major part of the motivation of this work. Due to this specific focus, this approach should be regarded in the mindset of uncertainty quantification in the broad field of scientific computing, where uncertainties are propagated through dynamical systems by simulation and their impact is quantified in the results.
There are several dimensions to further develop the application of the framework: a) more efficient numerical schemes for sampling the density functions in the Monte Carlo integration than, e.g. acceptance rejection sampling, b) more efficient use of the definition of the state space grid in the simulation, including possibly adaptive grids, c) better discretization approaches in the derivation of the random iteration equations (e.g. for continuous dynamical systems from Euler algorithm to Runge Kutta algorithms, e.g. [22]) and a detailed investigation of their convergence rates, d) extension of the optimization approaches, e.g. higher order iterative schemes such as full-density Newton optimization, etc., e) more challenging parameter density functions apart from normal distributions and a broader (also discontinuous) selection of the transfer functions, f) going from $R=2$ to higher dimensional state spaces and/or random variable vectors, e.g. investigating the curse of dimensionality for this approach, and finally, g) extension to other interesting application areas where iteration equations can play a productive role in understanding dynamical phenomena and contain random variable densities as a new degree of freedom in the modeling process.
5 Conclusion
In this paper, a unified framework for the simulation of random iteration equations (RIE) via direct probability density propagation has been developed. This is based on recent work on static random equations, which is iteratively applied in this study. In consequence, the algorithmic approach is conceptually efficient and straightforward. Expressive applications, such as the simulation of random and stochastic differential equations, the extension of gradient descent to the novel full-density gradient descent approach as well as the simulation of chaotic maps with strange attractors showcase its wide applicability. These demonstrations visualize complex stochastic evolutions of the simulation results and highlight the framework’s potential as a powerful tool for understanding the interplay between nonlinear dynamics and stochasticities in the field of uncertainty quantification. Several future directions of this approach are discussed.
6 Appendix
6.1 Numerical Verification with the 2D Ornstein-Uhlenbeck SDE
For an independent verification of the SDE calculation, we present a simple 2D Ornstein-Uhlenbeck SDE
$\text{d}x_{1}$ $=-x_{1}\,\text{d}t+\sigma_{1}\,\text{d}W_{1}$ (53)
$\text{d}x_{2}$ $=-x_{2}\,\text{d}t+\sigma_{2}\,\text{d}W_{2}\;,$ (54)
with the real diffusion parameters $\sigma_{1},\sigma_{2}$ and a Wiener process $\boldsymbol{W}_{t}$, leading to the transfer function
$\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})$ $:=\left(\begin{array}[]{c}x_{1}^{(n)}-x_{1}^{(n)}\,\Delta t+\sigma_{1}\,C_{1}^{(n)}\\ x_{2}^{(n)}-x_{2}^{(n)}\,\Delta t+\sigma_{2}\,C_{2}^{(n)}\\ \end{array}\right)\;.$ (57)
The stochastic noise densities are $f_{C_{1}},f_{C_{2}}:=\mathcal{N}(0,\Delta t)$ of the discritized SDE and the diffusion coefficients are $\sigma_{1}=0.4,\sigma_{2}=0.6$. In Figure 9 we utilize $P=384000$ samples, $f_{B_{1}},f_{B_{2}}:=\mathcal{N}(0,0.0025^{2})$ and $\Delta t=0.025$ and show the evolution of the density function. As starting distribution an (almost) point distribution at $\boldsymbol{x}_{0}=(1,0.8)^{T}$ is chosen (please note, we utilized a slightly extended square with size $0.045$ in order to utilize exactly the same algorithm as in Section 3.1). During this transient phase, the expected drift to the zero mean position is observable as well as the anisotropic increase in the diffusion.
The analytic solution for the mean $\boldsymbol{M}(t)$ and the covariance matrix $\boldsymbol{\Sigma}(t)$ for this process is known and given by
$\boldsymbol{M}(t)$ $=\boldsymbol{x}_{0}\,e^{-t}$ (58)
$\boldsymbol{\Sigma}(t)$ $=\frac{1}{2}\,(1-e^{-2\,t})\,\left(\begin{array}[]{cc}\sigma_{1}^{2}&0\\ 0&\sigma_{2}^{2}\end{array}\right)\;.$ (61)
In Figure 10 the numerical evaluation of the mean and the covariance values utilizing the full-density approach over 109 iterations in the transient phase are compared to the time-dependent analytic solution. The comparison demonstrates an agreement with minor differences verifying the calculation. Possible sources for remaining errors are the discretization errors (spatial as well as temporal utilizing the Euler-Maruyama approach), the finite number of Monte Carlo samples for integration, an extended starting distribution in the simulation instead of a point distribution as for the analytic solution and the slightly diffusing character of the standard deviation of $\boldsymbol{B}$ when solving the random iteration equations.
Figure 9: Illustration of the posterior densities for different iteration numbers of the 2D Ornstein-Uhlenbeck SDE model utilizing the full-density approach.
Figure 10: Illustration of the mean and covariance values empirically evaluated for the full-density algorithm (FD, solid lines) and the analytic solution (dashed lines). The strong similarity between FD and the analytic solution is demonstrated.

### 02 Apr 2026 v3

@@
 The presentation of the methodology is conceptually efficient based on recent work on static random equations and intentionally accessible.
-The technical requirements on the RIE are minimal based on the previous work, allowing for potential nonlinearities, discontinuities and stochasticities in the transfer function, as well as nonstandard densities and diffusion processes.
+The technical requirements on the RIE are mild based on the previous work, allowing for potential nonlinearities, discontinuities and stochasticities in the transfer function, as well as nonstandard densities and diffusion processes.
 As results, illustrative applications of random and stochastic differential equation simulations, a novel full-density gradient descent method (FDGD) for global optimization under uncertainty and examples of chaotic mappings are presented in order to demonstrate the breadth of the utility of this framework.
@@
 $\boldsymbol{x}^{(n+1)}-\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})=\boldsymbol{0}$ (2)
-leading to $\boldsymbol{B}$ a random variable vector centered around zero and with small standard deviations compared to other standard deviations and the required solution vector $\boldsymbol{x}^{(n+1)}$ for each iteration.
-The rather technical minimal assumptions about $\boldsymbol{T}$ are inherited directly from the integrability assumptions about $\boldsymbol{M}$ presented in [6] and are typically only violated in constructed limit cases.
+leading to $\boldsymbol{B}$ a random variable vector centered around zero and with small standard deviations compared to other standard deviations and the solution vector $\boldsymbol{x}^{(n+1)}$ for each iteration.
+The mild technical assumptions about $\boldsymbol{T}$ are inherited directly from the integrability assumptions about $\boldsymbol{M}$ presented in [6] and are typically only violated in constructed limit cases.
 In total, this leads to
@@
 Since $\boldsymbol{x}^{(n)}$ and $\boldsymbol{C}^{(n)}$ are independent, we get a likelihood distribution for $\boldsymbol{x}^{(n+1)}$ based on those densities and the specific transfer function $\boldsymbol{T}$ by
-$\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})=\int\limits_{\mathbb{R}^{K-R}}\int\limits_{\mathbb{R}^{R}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\;\text{d}\boldsymbol{s}_{1}\,\text{d}\boldsymbol{s}_{2}\;.$ (5)
+$\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})=\int\limits_{\mathbb{R}^{K}}\int\limits_{\mathbb{R}^{R}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\;\text{d}\boldsymbol{s}_{1}\,\text{d}\boldsymbol{s}_{2}\;.$ (5)
 which general form was presented previously [6].
@@
 We will leave such interesting possibilities outside of this study (although it is not outside of the framework) and only refer to the corresponding Section 2.5 in Hoegele [6].
+Rewriting Equation 5 into
+$\pi_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{x})\propto\int\limits_{\mathbb{R}^{R}}\underbrace{\left(\;\int\limits_{\mathbb{R}^{K}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\,\text{d}\boldsymbol{s}_{2}\right)}_{=:k(\boldsymbol{x},\boldsymbol{s}_{1})}\cdot\;\pi_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\;\text{d}\boldsymbol{s}_{1}\;,$ (7)
+the RIE iteration can be reframed as an integral linear operator acting on the posterior densities with kernel $k(\boldsymbol{x},\boldsymbol{s}_{1})$, which we denote as the linear RIE operator.
+We emphasize that this reformulation is in general possible including cases with nonlinearities, discontinuities or untypical parameter probability densities in the transfer function $\boldsymbol{T}$.
+Spectral investigation for specific kernels may lead to further insights into the dynamic properties of the density propagation for selected RIEs.
 2.2 Numerical Implementation
@@
 This is done in parallelization for speed-up and exactly $P$ samples are computed in each iteration.
+The major computational compromise must be found between the number of samples $P$, the standard deviations of $\boldsymbol{B}$ and the accepted noise level in the final density by utilizing Monte Carlo integration.
+For example, the more concentrated $\boldsymbol{B}$ is around zero (what is desirable since too large standard deviations introduce a blur on the posterior density or even slight diffusion over a large number of iterations), the more samples inserted in $f_{B_{r}}$ are typically not contributing to the final posterior density increasing the noise level.
 A possible schematic algorithm (which is also the basis of the presented results) looks like
@@
 We will present two of the most prominent use cases in the next subsections.
+It is pointed out that utilizing the perspective of the linear RIE operator combined with a linear decomposition of the posterior density function as linear combination of basis functions, much more efficient numerical schemes are possible than the presented Monte Carlo algorithm.
+In this case the coefficients of the linear combination are updated in each iteration by multiplying them with a propagation matrix.
+This perspective can also be utilized as an efficient method to incorporate spatial discretization directly into the algorithm by (effectively) pixel basis functions.
 2.3 Discretization of Dynamical Systems with Stochasticity as Random Iteration Equation
@@
 for an arbitrary dynamical system. Of course, more advanced discretization schemes can be utilized, but for simplicity we leave this demonstration applying the Euler algorithm.
-A major observation is that when utilizing this discretization scheme, we can also simulate the evolution of probability density functions of SDEs with random coefficients utilizing the Euler-Maruyama algorithm [4, 5, 10] by introducing new terms
+A major observation is that when utilizing this discretization scheme, we can also simulate the evolution of probability density functions of SDEs with random coefficients
+$\text{d}\boldsymbol{x}=\boldsymbol{F}(t,\boldsymbol{x},\boldsymbol{C})\,\text{d}t+\boldsymbol{B}(t,\boldsymbol{x},\boldsymbol{D})\,\text{d}\boldsymbol{W}_{t}\;,$ (12)
+with the drift coefficent function $\boldsymbol{F}$, the diffusion coefficient function $\boldsymbol{B}$ and a Wiener process $\boldsymbol{W}_{t}$, utilizing the Euler-Maruyama algorithm [4, 5, 10] by introducing new terms
 $\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}+\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)})\,\Delta t+\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{D}^{(n)})\,\Delta\boldsymbol{W}^{(n)}\;,$ (13)
-with the diffusion coefficient function $\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{D}^{(n)})$, which may depend on a random variable vector $\boldsymbol{D}^{(n)}$, and a Wiener process $\boldsymbol{W}_{t}$ with a discrete increment $\Delta\boldsymbol{W}^{(n)}=\boldsymbol{W}_{t_{n+1}}-\boldsymbol{W}_{t_{n}}$ as normally distributed random variables with zero mean and variances $\Delta t$.
-This leads to the general transfer function for SDEs
+with the coefficent random variable vectors $\boldsymbol{C}^{(n)}$ and $\boldsymbol{D}^{(n)}$, and the discrete Wiener process increment $\Delta\boldsymbol{W}^{(n)}=\boldsymbol{W}_{t_{n+1}}-\boldsymbol{W}_{t_{n}}$ as normally distributed random variables with zero mean and variances $\Delta t$.
+This leads to the general transfer function for SDEs by renaming the random variables
 $\boldsymbol{T}(\boldsymbol{x}^{(n)},\boldsymbol{C}^{(n)}):=\boldsymbol{x}^{(n)}+\boldsymbol{F}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}_{1..K_{1}}^{(n)})\,\Delta t+\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{C}_{K_{1}+1..K_{1}+K_{2}}^{(n)})\,\boldsymbol{C}_{K_{1}+K_{2}+1..K_{1}+K_{2}+R}^{(n)}\;,$ (14)
@@
 There are several dimensions to further develop the application of the framework: a) more efficient numerical schemes for sampling the density functions in the Monte Carlo integration than, e.g.
-acceptance rejection sampling, b) more efficient use of the definition of the state space grid in the simulation, including possibly adaptive grids, c) better discretization approaches in the derivation of the random iteration equations (e.g.
+acceptance rejection sampling, as well as utilizing the linear RIE operator combined with a linear decomposition of the posterior density functions, b) more efficient use of the definition of the state space grid in the simulation, including possibly adaptive grids, c) better temporal discretization approaches in the derivation of the random iteration equations (e.g.
 for continuous dynamical systems from Euler algorithm to Runge Kutta algorithms, e.g.